the number of legislators has increased approximately linearly from 1131 positions in 2002 to 1257 positions in 2009. Let n be the number of legislators at t years since 2000. Find an equation of a linear model to describe the data.
Answer by TimothyLamb(4379)

slope = m = dy/dx
m = (1257 - 1131)/(2009 - 2002)
m = 18
---
linear equation for the years 2002 to 2009:
n = 18t + 1131
---
n for the year 2000:
n = 18t + 1131
n = 18*(2000-2002) + 1131
n = 18*(-2) + 1131
n = 1095
---
answer:
linear equation for the years 2000 to 2009:
n = 18t + 1095
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php